# A Calculus of Overlays

### v2

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Overlay-calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are function color blind (Nystrom, 2015), ordinary arithmetic yields the relational semantics of logic programming, and self-reference resolves to multiple targets, making overlay-calculus strictly more expressive than the $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, in the spirit of F-coalgebras (Rutten, 2000), where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. Through iterative refinement, beginning with an early version implemented in Nix, the system evolved into the Overlay language (Yang, 2025b), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in the Overlay language, including boolean logic and natural number arithmetic (Section 5) and a web application (Section 5), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate overlay files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (Section 5); and that the same web application overlay file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section 5).
@@
This paper distills these observations into overlay-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into overlay-calculus, overlay-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into overlay-calculus requires five local, syntax-directed rules (Section 4), but the reverse is impossible—overlay enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
@@
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
@@
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. [^1] The Overlay language is statically typed and therefore stricter than the untyped overlay-calculus. All examples in this paper are ported from the Overlay language and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
@@
  • We present overlay-calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two overlays that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
@@
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into overlay-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section 5).
  • Open extension via overlay changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express overlay (Section 4.2).
@@ 2. Syntax
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in overlay-calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); overlay-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@
  • Named form: $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
@@
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
@@ 3. Overlay Trees
Unlike the $\lambda$-calculus, which requires reduction rules ($\beta$-reduction) to specify how terms compute, overlay-calculus has no reduction steps. There is no rewriting of terms, no evaluation order, and no notion of a “result” that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics. This is the observational perspective introduced in Section 1: like an F-coalgebra (Rutten, 2000), an overlay tree’s behavior is determined by observation (property presence), not by reduction.
Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes. One function has no object-oriented analogue: $\mathtt{overlays}$ collects the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
@@
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\;\bigr\}$
@@
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the identity-sharing paths $\mathtt{overlays}$ of $p$ itself, plus the $\mathtt{overlays}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site overlay back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
Overlays
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overlays}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
@@
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
@@
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
@@
The definitions above constitute the complete semantics of overlay-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In overlay-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.

### v4

# A Calculus of Inheritance

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Inheritance-calculus is distilled from MIXINv2, a practical implementation in which we observed further emergent phenomena: the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in MIXINv2, including boolean logic and natural number arithmetic (Section 5) and a web application (Section Function color blindness: ), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5); and that the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section Function color blindness: ).
@@
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the reverse is not macro-expressible under the equivalence notions fixed in Section 4.2 and Appendix A— inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
@@
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
@@
All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example. [^1] MIXINv2 is statically typed and therefore stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
@@
  • We present inheritance-calculus, a minimal computational model for declarative programming that contains only three constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two mixins that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in prior mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
@@
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to an inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; this semantics does not introduce a route-priority disambiguation rule.
  • The $\lambda$-calculus (in A-normal form) translates into inheritance-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence under $\lambda$-definable contexts (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
  • Open extension via inheritance changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express inheritance-calculus (Section 4.2).
@@ 2. Syntax
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in inheritance-calculus. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@
Named form:
$\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
Indexed form:
$\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
@@
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that mixin, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
@@ 3. Mixin Trees
A mixin tree has a single abstraction, the mixin: every mixin can serve as field, method, class, and object at the same time in object-oriented terms.
Unlike the $\lambda$-calculus, which requires $\beta$-reduction to specify how terms compute, inheritance-calculus has no reduction steps: no rewriting of terms, no evaluation order, and no notion of a result that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
Each function has an analogue in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, $\mathtt{bases}$ collects the direct base classes, and $\mathtt{overrides}$ generalizes method override to the tree setting, collecting the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
@@
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\;\bigr\}$
@@
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{override}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\;\bigr\}$
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: $\mathtt{overrides}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overrides}$ first.
Overrides
Because inheritance-calculus has no separate notion of “method,” a mixin simultaneously plays the role of a class and the role of a method in object-oriented languages. When a mixin is viewed as a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label—this is method override. Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overrides}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
@@
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overrides}$ one step:
@@
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose override component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{override}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
@@
The definitions above constitute the complete semantics of inheritance-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In inheritance-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. Identity is determined by observable behavior rather than by internal structure.